OpenStudy (anonymous):

Find an exact value. tan 75°

OpenStudy (campbell_st):

use the sum of 2 angles tan(75) = tan(45 + tan 30) \[\tan(A + B) = \frac{\tan(A) + \tan(B)}{1 - \tan(A)\tan(B)}\] the exact values for tan(45) and tan(30) can be obtained from the exact value triangles